# Dynamic Subgraph Matching via Cost-Model-based Vertex Dominance Embeddings (Technical Report)

### v2

@@ 4.1. Preliminaries and Terminologies
As an example in Figure 2(a), the unit star subgraph $g_{v_{1}}$ in data graph $G_{t}$ contains a center vertex $v_{1}$ and its 1-hop neighbors $v_{2}$, $v_{3}$, and $v_{4}$. Figure 2(b) illustrates 8 ($=2^{3}$) possible star substructures $s_{v_{1}}$ ($\subseteq g_{v_{1}}$), which are centered at vertex $v_{1}$ and with different combinations of $v_{1}$’s 1-hop neighbors.
@@ 4.2. Vertex Dominance Embedding
In this subsection, we discuss how to obtain an embedding vector, $o(v_{i})$, for each vertex $v_{i}$, which embeds labels of vertex $v_{i}$ and its 1-hop neighbors (i.e., unit star subgraph or star substructure). Specifically, the embedding vector $o(v_{i})$ consists of two portions, a Seeded PseUdo Random (SPUR) vector $x_{i}$ and a SPUR Aggregated Neighbor (SPAN) vector $y_{i}$.
(a) unit star subgraph $g_{v_{1}}$
(b) 8 possible star substructures $s_{v_{1}}$
Figure 2. Illustration of unit star subgraph and star substructures.
@@
Seeded PseUdo Random (SPUR) Vector, $x_{i}$: In the dynamic graph $G_{D}$, each vertex $v_{i}$ is associated with a label $l(v_{i})$, which can be encoded by a nonnegative integer. By treating the vertex label $l(v_{i})$ as the seed, we can generate a Seeded PseUdo Random (SPUR) vector, $x_{i}$, of arity $d$ via a pseudo-random number generator $f_{1}(\cdot)$. That is, we have:
@@
SPUR Aggregated Neighbor (SPAN) Vector, $y_{i}$: Next, for each vertex $v_{i}$, we consider its 1-hop neighbors, $v_{j}\in\mathcal{N}_{v_{i}}$, in a star subgraph pattern (i.e., the unit star subgraph $g_{v_{i}}$ or star substructure $s_{v_{i}}$), and sum up their corresponding SPUR vectors $x_{j}$ to generate a SPur Aggregated Neighbor (SPAN) vector $y_{i}$:
@@
where ||||| | is the concatenation operation of two vectors, and $x_{i}$ and $y_{i}$ are given in Eqs. (1) and (2), respectively.
Intuitively, vertex dominance embedding $o(v_{i})$ encodes label features of a star subgraph pattern (i.e., $g_{v_{i}}$ or $s_{v_{i}}$) containing center vertex $v_{i}$ and its 1-hop neighbors. To distinguish vertex dominance embeddings $o(v_{i})$ from different star subsgraph patterns, we also use notations $o(g_{v_{i}})$ or $o(s_{v_{i}})$.
@@ 4.3. Properties of Vertex Dominance Embedding
The Dominance Property of Vertex Embeddings: From Eq. (3), the vertex dominance embedding $o(v_{i})$ is a combination of SPUR and SPAN vectors. Given any unit star subgraph $g_{v_{i}}$ and its star substructure $s_{v_{i}}$ (i.e., $s_{v_{i}}\subseteq g_{v_{i}}$), their vertex dominance embeddings always satisfy the dominance (Borzsony et al., 2001) (or equality) condition: $o(s_{v_{i}})[j]\leq o(g_{v_{i}})[j]$, for all dimensions $1\leq j\leq 2d$, which is denoted as $o(s_{v_{i}})\preceq o(g_{v_{i}})$ (including the case where $o(s_{v_{i}})=o(g_{v_{i}})$).
@@
(The Dominance Property of Vertex Embeddings) Given a unit star subgraph $g_{v_{i}}$ centered at vertex $v_{i}$ and any of its star substructures $s_{v_{i}}$ (i.e., $s_{v_{i}}\subseteq g_{v_{i}}$), their vertex dominance embeddings satisfy the condition that: $o(s_{v_{i}})\preceq o(g_{v_{i}})$ (including $o(s_{v_{i}})=o(g_{v_{i}})$) in the embedding space.
@@
The Usage of the Vertex Embedding Property: Note that, a star substructure $s_{v_{i}}$ ($\subseteq g_{v_{i}}$) can be a potential query unit star subgraph $s_{q_{i}}$ from the query graph $q$ (containing a center query vertex $q_{i}$ and its 1-hop neighbors) for dynamic subgraph matching.
During the subgraph matching, given a query embedding vector $o(q_{i})$ of query vertex $q_{i}$, we can identify those candidate matching vertices $v_{i}$ in dynamic graph $G_{D}$ with embeddings $o(v_{i})$ dominated by $o(q_{i})$. This way, we can convert the dynamic subgraph matching problem into a dominance search problem in the embedding space.
Figure 4. An example of vertex dominance embeddings in subgraph matching ( ||||| | is the concatenation operation of two vectors).
@@
Ease of Incremental Updates for Vertex Dominance Embeddings: In the dynamic graph $G_{D}$, our proposed vertex dominance embedding $o(v_{i})$ can be incrementally maintained, upon graph update operations in $\Delta G_{t}$.
Specifically, any edge insertion $(e+,t)$ or deletion $(e-,t)$ (as given in Definition 2.4) will affect the embeddings of two ending vertices, $v_{i}$ and $v_{j}$, of edge $e=(v_{i},v_{j})$. We consider incremental update of the embedding vector $o(v_{i})$ for vertex $v_{i}$ as follows.
@@
The embedding update of vertex $v_{j}$ is similar and thus omitted.
Note that, for each graph update (i.e., edge insertion $(e+,t)$ or deletion $(e-,t)$), the time complexity of incrementally updating vertex dominance embedding vectors is given by $O(d)$, where $d$ is the dimension of SPUR/SPAN vectors.
Figure 5. An example of vertex embeddings for DSM.
@@
Figure 5 illustrates an example of using our vertex dominance embeddings to conduct the subgraph matching over dynamic graph $G_{D}$ over time $t$ from 0 to 3, where $q$ is a given query graph. Consider two vertices $v_{1}$ and $q_{1}$ in graphs $G_{D}$ and $q$, respectively. At timestamp $t=0$, their vertex embeddings satisfy the dominance relationship, that is, $o(q_{1})\preceq o(v_{1})$, thus, vertex $v_{1}$ is a candidate matching with $q_{1}$.
@@
(a) SPUR vectors $x_{i}$
(b) SPAN vectors $y_{i}$
Figure 6. An example of SPUR/SPAN vector distributions in vertex dominance embedding $o(v_{i})$.
@@ 4.4. Embedding Optimization with Base Vector $z_{i}$
Figure 6 shows the distributions of 2D SPUR and SPAN vectors, $x_{i}$ and $y_{i}$, respectively, in vertex dominance embedding $o(v_{i})$ over Yeast graph (Sun and Luo, 2020) (with 3,112 vertices and 12,519 edges). We can see that, in Figure 6(a), the SPUR vectors $x_{i}$ generated by a seeded randomized function are distributed uniformly in the embedding space, whereas the SPAN vectors (i.e., the summed SPUR vectors) $y_{i}$ are more clustered along the reverse diagonal line in Figure 6(a).
As mentioned in Section 4.3, given a query vertex $q_{i}$, we can use its embedding vector $o(q_{i})$ as a query point to find the dominated embedding vectors $o(v_{i})$ ($=x_{i}\text{ }||\text{ }y_{i}$) in the 4D embedding space. Due to the scattered SPUR/SPAN vectors in the embedding space, it is very likely that some false alarms of embedding vectors $o(v_{i})$ are included as candidate matching vertices (i.e., dominated by $o(q_{i})$) .
In order to further enhance the pruning power, our goal is to update our vertex dominance embedding vector from $o(v_{i})$ to $o^{\prime}(v_{i})$, such that the number of candidate vertices dominated by $o(q_{i})$ is reduced (i.e., as small as possible). Specifically, we have the optimized embedding $o^{\prime}(v_{i})$, given by:
@@
Note that, since the base vector $z_{i}$ is generated only based on the label $l(v_{i})$ of center vertex $v_{i}$, different star subgraphs with the same center vertex label $l(v_{i})$ will result in the same base vector $z_{i}$. Therefore, the new (optimized) vertex embedding vector $o^{\prime}(v_{i})$ in Eq. (4) still follows the dominance relationship, that is, $o^{\prime}(q_{i})\preceq o^{\prime}(v_{i})$ holds, if a query star pattern $s_{q_{i}}$ is a subgraph of unit star subgraph $g_{v_{i}}$ (i.e., $s_{q_{i}}\subseteq g_{v_{i}}$ and $l(q_{i})\equiv l(v_{i})$).
@@
(The Dominance Property of the Optimized Vertex Embeddings) Given a unit star subgraph $g_{v_{i}}$ centered at vertex $v_{i}$ and any of its star substructures $s_{v_{i}}$ (i.e., $s_{v_{i}}\subseteq g_{v_{i}}$), their optimized vertex embeddings (via base vector $z_{i}$) satisfy the condition that: $o^{\prime}(s_{v_{i}})\preceq o^{\prime}(g_{v_{i}})$ (including $o^{\prime}(s_{v_{i}})=o^{\prime}(g_{v_{i}})$).
@@
Due to the property of vertex dominance embeddings (as given in Lemma LABEL:lemma:embedding_property), we have $o(s_{v_{i}})\preceq o(g_{v_{i}})$, or equivalently $\alpha o(s_{v_{i}})\preceq\alpha o(g_{v_{i}})$ (for $\alpha>0$).
@@
Discussions on How to Design a Base Vector, $z_{i}$: To improve the pruning power, we would like to make the distribution of the updated vertex embeddings in Eq. (4) more dispersed along diagonal line (or plane), so that fewer false alarms (dominated by $o^{\prime}(q_{i})$) can be obtained during dynamic subgraph matching.
In this paper, we propose to design the base vector $z_{i}$ as a random vector distributed on the unit diagonal line/hyperplane ($L_{1}$-norm) in the first quadrant. That is, we can normalize $z_{i}$ to $\frac{z_{i}}{||z_{i}||_{1}}$, where $||\cdot||_{1}$ is $L_{1}$-norm. Moreover, since constant $\alpha$ is far smaller than $\beta$, $\alpha(x_{i}||y_{i})$ can be considered as noises added to the scaled base vector $\beta z_{i}$. Thus, vertex embedding vectors $o^{\prime}(v_{i})$ ($=\alpha(x_{i}||y_{i})+\beta z_{i}$) in Eq. (4) are still distributed close to a diagonal line (or hyperplane) in the first quadrant.
@@
Figure 7 shows the distributions of the updated 2D SPUR and SPAN vectors (from $o^{\prime}(v_{i})$), by adding a base vector $z_{i}$, where $z_{i}$ is given by $L_{1}$-norm. Since we have $\alpha\ll\beta$ in Eq. (4), after adding $\alpha(x_{i}||y_{i})$ to the scaled base vector $\beta z_{i}$, the previous embedding vector $o(v_{i})$ (as shown in Figure 6) is re-located close to diagonal lines in Figure 7. This way, we can reduce the number of vertex false alarms dominated by $o(q_{i})$ in both SPUR and SPAN spaces.
(a) $\alpha x_{i}+\beta(z_{i}[0],z_{i}[1])^{T}$
(b) $\alpha y_{i}+\beta(z_{i}[2],z_{i}[3])^{T}$
Figure 7. An example of $o^{\prime}(v_{i})$ distributions with $L_{1}$-norm diagonal-line base vector $z_{i}$ ($\alpha=0.01,\beta=100$).
@@ 5. Degree-Aware Star Substructure Synopses
In this section, we will build effective synopses over embeddings of star substructures of different degrees in dynamic graph $G_{D}$, which can be incrementally updated and support efficient DSM query processing.
@@ 5.1. Vertex Embedding via Degree Grouping
Note that, for the subgraph matching over a dynamic graph $G_{t}$, those vertices $v_{i}\in V(G_{t})$ with high degrees $deg(v_{i})$ are less likely to be pruned in the embedding space. This is because, for high degrees, the SPAN vectors in $o^{\prime}(v_{i})$ tend to have large values. Thus, $o^{\prime}(v_{i})$ is more likely to be dominated by query embedding vector $o^{\prime}(q_{i})$, and treated as candidate vertices for the refinement (i.e., $v_{i}$ cannot be pruned).
To enhance the pruning power for high-degree vertices, in this subsection, we propose a novel and effective degree grouping technique, which separately maintains vertex embeddings, $o^{\prime}(v_{i})$, for star substructures $s_{v_{i}}$ of different (lower) degree groups (instead of unit star subgraph $g_{v_{i}}$) in $G_{t}$.
Equi-Frequency Degree Grouping: One straightforward way to perform the degree grouping is to let each single degree value be a degree group. However, since vertex degrees in real-world graphs usually follow the power-law distribution (Barabási and Albert, 1999; Newman, 2005; Barabási and Bonabeau, 2003; Albert and Barabási, 2002; Dorogovtsev and Mendes, 2003), only a small fraction of vertices have high degrees. In other words, only a few unit star subgraphs $g_{v_{i}}$ have star substructures $s_{v_{i}}$ of very high degrees, which is not space- and time-efficient to maintain in a dynamic graph $G_{t}$.
Inspired by this, in this paper, we propose to maintain vertex embeddings for star substructures with $m$ degree groups (of different degree intervals). We assume that the degree statistics of the initial graph $G_{0}$ are similar to that in $G_{t}$.
This way, we obtain the probability mass function of vertex degrees in $G_{0}$ is $pmf(\delta)=\frac{freq(deg(v_{i})\geq\delta)}{|V(G_{0})|}$, where $freq(deg(v_{i})\geq\delta)$ is the number of vertices $v_{i}\in V(G_{0})$ whose degrees are greater than or equal to a degree value $\delta$. Our goal is to divide the degree interval, $[1,deg_{max}]$, into $m$ degree groups =$\mathcal{B}=\{B_{1},B_{2},\cdots,B_{m}\}$, such that each degree bucket $B_{j}$ ($=(\delta_{j-1},\delta_{j}]$) contains the same (or similar) mass in the distribution of $pmf(\delta)$ (for $\delta\in B_{j}$).

### v3

@@ 4.1. Preliminaries and Terminologies
As an example in Figure LABEL:subfig:subgraph, the unit star subgraph $g_{v_{1}}$ in data graph $G_{t}$ contains a center vertex $v_{1}$ and its 1-hop neighbors $v_{2}$, $v_{3}$, and $v_{4}$. Figure LABEL:subfig:substructure shows 8 ($2^{3}$) possible star substructures $s_{v_{1}}$ ($\subseteq g_{v_{1}}$), which are centered at vertex $v_{1}$ and with different combinations of $v_{1}$’s 1-hop neighbors.
Note that, intuitively, star substructures $s_{v_{i}}$ are potential query star patterns (i.e., vertex $q_{i}\in V(q)$ and its 1-hop neighbors) in the query graph $q$ of the CSM problem.
@@ 4.2. Vertex Dominance Embedding
To tackle the CSM problem, our goal is to avoid costly graph comparisons by transforming them into a search problem over vertex embeddings in a vector space. Thus, we design an embedding method to generate a vector, $o(v_{i})$, for each vertex $v_{i}\in V(G_{t})$ based on the labels of vertex $v_{i}$ and its 1-hop neighbors (i.e., unit star subgraph or star substructure). Specifically, the embedding vector $o(v_{i})$ consists of two portions, a Seeded PseUdo Random (SPUR) vector $x_{i}$ and a SPUR Aggregated Neighbor (SPAN) vector $y_{i}$.
(a) Illustration of unit star subgraph and substructures.
@@
Seeded PseUdo Random (SPUR) Vector, $x_{i}$: In the dynamic graph $G_{D}$, each vertex $v_{i}$ is associated with a label $l(v_{i})$, which can be encoded by a nonnegative integer. By treating the vertex label $l(v_{i})$ as the seed, we can generate a Seeded PseUdo Random (SPUR) vector, $x_{i}$, of a given tunable arity $d$ via a pseudo-random number generator $f_{1}(\cdot)$, that is,
@@
SPUR Aggregated Neighbor (SPAN) Vector, $y_{i}$: For each vertex $v_{i}$, we also aggregate structural information from its 1-hop neighbors, $v_{j}\in\mathcal{N}_{v_{i}}$, in a star subgraph pattern (i.e., unit star subgraph $g_{v_{i}}$ or star substructure $s_{v_{i}}$), and sum up their corresponding SPUR vectors $x_{j}$ to obtain a SPur Aggregated Neighbor (SPAN) vector $y_{i}$:
@@
where $||$ is the concatenation operation of two vectors, and $x_{i}$ and $y_{i}$ are given in Eqs. (1) and (2), respectively.
Intuitively, vertex dominance embedding $o(v_{i})$ encodes label features of a star subgraph pattern (i.e., $g_{v_{i}}$ or $s_{v_{i}}$) containing center vertex $v_{i}$ and its 1-hop neighbors. To distinguish vertex dominance embeddings $o(v_{i})$ from different star subgraph patterns, we also use notations $o(g_{v_{i}})$ or $o(s_{v_{i}})$.
@@
Figure 5. An example of vertex dominance embeddings in subgraph matching ($||$ is the concatenation of two vectors).
@@ 4.3. Properties of Vertex Dominance Embedding
The Dominance Property of Vertex Embeddings: From Eq. (3), the vertex embedding $o(v_{i})$ is a combination of SPUR and SPAN vectors. Given any unit star subgraph $\bm{g_{v_{i}}}$ and its star substructure $\bm{s_{v_{i}}}$ (i.e., $\bm{s_{v_{i}}\subseteq g_{v_{i}}}$), their vertex dominance embeddings always satisfy the dominance (Borzsony et al., 2001) (or equality) condition: $\bm{o(s_{v_{i}})[j]\leq o(g_{v_{i}})[j]}$, for all dimensions $\bm{1\leq j\leq 2d}$, denoted as $o(s_{v_{i}})\preceq o(g_{v_{i}})$ (including the case that $o(s_{v_{i}})=o(g_{v_{i}})$). Here, our (relaxed) dominance semantics include both classic dominance (in skyline definition (Borzsony et al., 2001)) and equality cases. Formally, for two embedding vectors $o(x),o(y)\in\mathbb{R}^{2d}$, we say that: $o(x)$ dominates $o(y)$, if $o(x)[i]\leq o(y)[i]$ holds for all dimensions $1\leq i\leq 2d$.
@@
(The Dominance Property of Vertex Embeddings) Given a unit star subgraph $g_{v_{i}}$ centered at vertex $v_{i}$ and any of its star substructures $s_{v_{i}}$ (i.e., $s_{v_{i}}\subseteq g_{v_{i}}$), their vertex embeddings satisfy the dominance condition that: $o(s_{v_{i}})\preceq o(g_{v_{i}})$ (including $o(s_{v_{i}})=o(g_{v_{i}})$) in the embedding space.
@@
The Usage of the Vertex Embedding Property: Note that, a star substructure $s_{v_{i}}$ ($\subseteq g_{v_{i}}$) can be a potential query unit star subgraph $s_{q_{i}}$ from the query graph $q$ (containing a center query vertex $q_{i}$ and its 1-hop neighbors) for continuous subgraph matching.
During the subgraph matching, given a query embedding vector $o(q_{i})$ of query vertex $q_{i}$, we can identify candidate data vertices $v_{i}$ in dynamic graph $G_{D}$ with embeddings $o(v_{i})$ dominated by $o(q_{i})$. This way, we can convert the continuous subgraph matching problem into a dominance search problem in the embedding space.
@@
Figure 6. An example of vertex embeddings for CSM.
Ease of Incremental Updates for Vertex Dominance Embeddings: In the dynamic graph $G_{D}$, our proposed vertex dominance embedding $o(v_{i})$ can be incrementally maintained, upon graph update operations in $\Delta G_{t}$. Specifically, any edge insertion $(e+,t)$ or deletion (e,t)(e-,t) operation (as given in Definition 2.4) will affect the embeddings of two ending vertices, $v_{i}$ and $v_{j}$, of edge $e=(v_{i},v_{j})$. We incrementally update the embedding $o(v_{i})$ for vertex $v_{i}$ below.
@@
The case of updating vertex embedding $o(v_{j})$ is similar and omitted.
Note that, for each graph update (i.e., edge insertion $(e+,t)$ or deletion (e,t)(e-,t)), the time complexity of incrementally updating vertex dominance embedding vectors is given by $O(d)$, where $d$ is the dimension of SPUR/SPAN vectors.
@@
Figure 6 illustrates an example of using our vertex dominance embeddings to conduct CSM over the dynamic graph $G_{D}$ over time $t$ from 0 to 3, where $q$ is a given query graph. Consider two vertices $v_{1}$ and $q_{1}$ in graphs $G_{D}$ and $q$, respectively. At timestamp $t=0$, their embeddings satisfy the dominance relationship, that is, $o(q_{1})\preceq o(v_{1})$, thus, vertex $v_{1}$ is a candidate matching with $q_{1}$.
@@ 4.4. Embedding Optimization with Base Vector $z_{i}$
Figure 8(a) shows the distributions of 2D SPUR and SPAN vectors, $x_{i}$ and $y_{i}$, respectively, in vertex dominance embedding $o(v_{i})$ over Yeast graph (Sun and Luo, 2020) (with 3,112 vertices and 12,519 edges). We can see that, in Figure LABEL:subfig:spur_vectors, the SPUR vectors $x_{i}$ generated by a seeded randomized function are distributed uniformly in the embedding space, whereas the SPAN vectors (i.e., the summed SPUR vectors) $y_{i}$ are more clustered along the reverse diagonal line in Figure LABEL:subfig:span_vectors.
As mentioned in Section 4.3, given a query vertex $q_{i}$, we can use its embedding vector $o(q_{i})$ as a query point to find the dominated embedding vectors $o(v_{i})$ ($x_{i}\text{ }||\text{ }y_{i}$) in the 4D embedding space. Due to the scattered SPUR/SPAN vectors in the embedding space, it is very likely that some false alarms of embedding vectors $o(v_{i})$ are included as candidate matching vertices (i.e., dominated by $o(q_{i})$).
(a) An example of SPUR/SPAN vector distributions in vertex dominance embedding $o(v_{i})$.
(b) An example of $o^{\prime}(v_{i})$ distributions with $L_{1}$-norm diagonal-line base vector $z_{i}$ ($\alpha=0.01$ and $\beta=100$).
In order to further enhance the pruning power, our goal is to update our vertex dominance embedding vector from $o(v_{i})$ to $o^{\prime}(v_{i})$, such that the number of candidate vertices dominated by $o(q_{i})$ is reduced (i.e., as small as possible). Specifically, we propose to use an optimized embedding $o^{\prime}(v_{i})$, given by:
@@
Note that, since the base vector $z_{i}$ is generated only based on the (seed) label $l(v_{i})$ of center vertex $v_{i}$, different star subgraphs with the same center vertex label $l(v_{i})$ will result in the same base vector $z_{i}$. Therefore, the new (optimized) vertex embedding vector $o^{\prime}(v_{i})$ in Eq. (4) still follows the dominance relationship, that is, $o^{\prime}(q_{i})\preceq o^{\prime}(v_{i})$ holds, if a query star pattern $s_{q_{i}}$ is a subgraph of unit star subgraph $g_{v_{i}}$ (i.e., $s_{q_{i}}\subseteq g_{v_{i}}$ and $l(q_{i})\equiv l(v_{i})$).
@@
(The Dominance Property of the Optimized Vertex Dominance Embeddings) Given a unit star subgraph $g_{v_{i}}$ centered at vertex $v_{i}$ and any of its star substructures $s_{v_{i}}$ (i.e., $s_{v_{i}}\subseteq g_{v_{i}}$), their optimized vertex dominance embeddings (via base vector $z_{i}$) satisfy the condition that: $o^{\prime}(s_{v_{i}})\preceq o^{\prime}(g_{v_{i}})$ (including $o^{\prime}(s_{v_{i}})=o^{\prime}(g_{v_{i}})$) in the embedding space.
@@
Due to the property of vertex dominance embeddings (as given in Lemma 4.1), we have $o(s_{v_{i}})\preceq o(g_{v_{i}})$, or equivalently $\alpha o(s_{v_{i}})\preceq\alpha o(g_{v_{i}})$ (for $\alpha>0$).
@@
Figure 8. An example of degree grouping and degree-aware star substructure synopses ($\alpha=0.01$, $\beta=100$, and $z_{1}=(0.36,0.64)$).
Discussions on How to Design a Base Vector, $z_{i}$: To improve the pruning power, we would like to make the distribution of the updated vertex embeddings in Eq. (4) more dispersed along diagonal line (or plane), so that fewer false alarms (dominated by $o^{\prime}(q_{i})$) can be obtained during continuous subgraph matching.
[…]
@@
Figure 8(b) shows the distributions of the updated 2D SPUR and SPAN vectors (from $o^{\prime}(v_{i})$), by adding a base vector $z_{i}$, where $z_{i}$ is given by $L_{1}$-norm. Since we have $\alpha\ll\beta$ in Eq. (4), after adding $\alpha(x_{i}||y_{i})$ to the scaled base vector $\beta z_{i}$, the previous embedding vector $o(v_{i})$ (as shown in Figure 8(a)) is re-located close to diagonal lines in Figure 8(b). This way, we can reduce the number of vertex false alarms dominated by $o(q_{i})$ in both SPUR and SPAN spaces, and in turn enhance the CSM pruning power.
@@ 5. Degree-Aware Star Substructure Synopses
We observed that high-degree center vertices of star substructures tend to have vertex dominance embeddings with low pruning power during CSM processing. Therefore, in this section, we will classify star substructures into different degree groups, and build synopses over their embeddings under degree groups. This way, CSM can be conducted over a synopsis corresponding to a group of lower degrees, which can further improve the pruning power.
@@ 5.1. Vertex Embedding via Degree Grouping
Rationale behind the Degree Grouping: During the subgraph matching over a dynamic graph $G_{t}$, those vertices $v_{i}\in V(G_{t})$ with high degrees $deg(v_{i})$ are less likely to be pruned in the embedding space. This is because, for high degrees, the SPAN vectors in $o^{\prime}(v_{i})$ tend to have large values on all dimensions. Thus, $o^{\prime}(v_{i})$ also tends to be dominated by query embedding vector $o^{\prime}(q_{i})$, and treated as candidate vertices for the refinement (i.e., $v_{i}$ cannot be pruned).
Therefore, to enhance the pruning power for high-degree vertices, in this subsection, we propose a novel and effective degree grouping technique, which divides all possible star substructures $s_{v_{i}}$ of each vertex $v_{i}$ in $G_{D}$ (that may match with a query star pattern $s_{q_{i}}$) into different groups (corresponding to different degree intervals of center vertices $v_{i}$ in star substructures $s_{v_{i}}$). Intuitively, during the CSM query processing, the embedding vector $o^{\prime}(q_{i})$ of query star pattern $s_{q_{i}}$ only needs to be compared with those of star substructures $s_{v_{i}}$ with similar degrees in a group, rather than all star substructures (including those with high degrees), which can greatly improve the pruning power.
Equi-Frequency Degree Grouping: One straightforward way to perform the degree grouping is to let each single degree value be a degree group. However, since vertex degrees in real graphs usually follow the power-law distribution (Barabási and Albert, 1999; Newman, 2005; Barabási and Bonabeau, 2003; Albert and Barabási, 2002; Dorogovtsev and Mendes, 2003), only a small fraction of vertices have high degrees. That is, only a few unit star subgraphs $g_{v_{i}}$ have star substructures $s_{v_{i}}$ of very high degrees, which is not space-/time-efficient to maintain in the dynamic graph $G_{t}$.
Inspired by this, in this paper, we propose to maintain vertex embeddings for star substructures with $m$ degree groups (of different degree intervals). We assume that the degree statistics of the initial graph $G_{0}$ are similar to those in $G_{t}$.
This way, we obtain the probability mass function of vertex degrees in $G_{0}$ is $pmf(\delta)=\frac{freq(deg(v_{i})\geq\delta)}{|V(G_{0})|}$, where $freq(deg(v_{i})\geq\delta)$ is the number of vertices $v_{i}\in V(G_{0})$ whose degrees are greater than or equal to a degree value $\delta$. Our goal is to divide the degree interval, $(0,deg_{max}]$, into $m$ degree groups $\mathcal{B}=\{B_{1},B_{2},\cdots,B_{m}\}$, such that each degree bucket $B_{j}$ ($(\delta_{j-1},\delta_{j}]$) contains the same (or similar) mass in the distribution of $pmf(\delta)$ (for $\delta\in B_{j}$).
Example 0.
Figure 8 illustrates the degree grouping process for star substructures with center vertex $v_{1}$, where 7 possible star substructures $s_{v_{1}}^{i}$ (excluding Case ⑦, the isolated vertex $v_{1}$) are partitioned into two groups with 2 degree intervals, $B_{1}=(0,1]$ and $B_{2}=(1,3]$, respectively. For example, in Case ①, center vertex $v_{1}$ in star substructure $s_{v_{1}}^{1}$ has degree 1, which thus falls into the first degree bucket $B_{1}=(0,1]$.
A query star pattern $s_{q_{1}}$ with center vertex $q_{1}$ of degree 1 will be compared with star substructures in degree bucket $B_{1}=(0,1]$ only. In contrast, for a query star pattern $s_{q_{1}}$ with degree equal to 2 or 3, we only need to search the second degree bucket $B_{2}$ (with interval $(1,3]$).